OpenStudy (anonymous):

Trains A and B are traveling in the same direction on parallel tracks. Train A is traveling at 40 miles per hour and train B is traveling at 48 miles per hour. Train A passes a station at 4:10 a.m. If train B passes the same station at 4:25 a.m., at what time will train B catch up to train A?

OpenStudy (anonymous):

Here's how I would solve this: 1) Craft functions f(x) and g(x) for positions of the trains 2) Solve f(x) = g(x). Both functions are linear, so there is only one solution.

OpenStudy (anonymous):

Simultaneous equations might be a better idea though. :)

OpenStudy (anonymous):

let t=elapsed time for B to catchup with A. Solve the following for t:\[40 \left(t+\frac{15}{60}\right)=48 t \]

OpenStudy (anonymous):

they r asking for a time?????

OpenStudy (anonymous):

The intercept time of day is 4:25AM + t or 4:25 + (5/4 hours converted to 1:15) = 4:25 +1:15 = 5:35 AM. But, I could be wrong.

OpenStudy (anonymous):

how is the conversion done from 5/4 to 1:15???

OpenStudy (anonymous):

it would be 5:40 not 35

OpenStudy (anonymous):

5/4=1+1/4, A quarter of an hour is 15 minutes. Another way to do the calculation is to multiply 1/4 by 60 minutes. 5/4 hours is 1 hour plus 15 minutes or 1:15 in the standard hour/minute notation. Yes, it appears that I cannot add 25 minutes to 15 minutes and get the right answer this morning. Thank you for the alert. Wouldn't it be nice if we had some kind of metric hour to replace the normal hour.